1. A ladder 20 ft long leans against a vertical building. If the top of the ladder slides down at a rate of p 3 ft s, how fast is the bottom of the ladder sliding away from the building when the top of the ladder is 10 ft above the ground?



2. A sphere is growing in such a manner that its volume increases at 0:2 mΒ³ s(cubic meter per second). How fast is its radius increasing when it is 7 m long?

1
Expert's answer
2022-03-05T05:04:42-0500

1.


"x^2+y^2=(20)^2"

Differentiate both sides with respect to "t"


"2x(dx\/dt)+2y(dy\/dt)=0"

"dx/dt=-(\dfrac{y}{x})dy/dt"

"dx/dt=-(\dfrac{y}{\sqrt{400-y^2}})dy/dt"

Given "dy\/dt=-3ft\/s, y=10ft"


"dx/dt=-(\dfrac{10}{\sqrt{400-(10)^2}})(-3ft/s)"

"dx/dt=\sqrt{3}\ ft/s"

2.

"V=\dfrac{4\pi}{3} r^3"

Differentiate with respect to "t"


"dV/dt=\dfrac{4\pi}{3}(3 r^2)dr/dt"

"dr/dt=\dfrac{1}{4\pi r^2}dV/dt"

"dr/dt=\dfrac{1}{4\pi (7m)^2}(0.2m^3/s)\approx0.000325\ m/s"